# From Bias to Balance: Exploring and Mitigating Spatial Bias in LVLMs

### v1

@@ 5.5 Analysis of Information Flow
To analyze the effect of BaPA on cross-modal interactions, we visualize the information flow in LLaVA-v1.6 and the BaPA-finetuned version across three well-performing downstream datasets, including ScienceQA, CRPE, and HallusionBench. For each dataset, we compute the average attention score from text tokens to image tokens across all layers and attention heads, thereby highlighting the contribution of each image token during multimodal reasoning.
Figure 5: The visualization of information flow from image token $\tilde{v_{i}}$ to text token $x_{j}$.
Overall Trend. The visualization of information flow across downstream datasets shown in Figure 5(a) reveals clear differences between LLaVA-v1.6 and the BaPA-finetuned version. The baseline model tends to concentrate attention on only a few image tokens, often overlooking other semantically relevant regions. This behavior indicates that its cross-modal interaction relies heavily on localized cues, which can increase the risk of biased or insufficient interactions. In contrast, BaPA encourages a more even spread of attention across image tokens, ensuring that a broader set of important visual features contributes to the cross-modal interaction. These results demonstrate that BaPA effectively stabilizes attention patterns, allowing LVLMs to capture richer and more comprehensive visual information during cross-modal interaction.
Case study. To better illustrate the impact of BaPA, we present a case study from HallusionBench in Figure 5(b). The baseline LLaVA-v1.6 confines its attention to a few isolated image regions, while most visual tokens receive little to no attention. This uneven distribution reflects the position imbalance introduced by standard position assignment, leading the model to overemphasize spurious cues while neglecting relevant evidence. By contrast, LLaVA-v1.6-BaPA exhibits a more balanced allocation of attention across image tokens, particularly in regions strongly associated with correct decision—making (i.e., tokens corresponding to the curve and words). This balanced allocation facilitates the integration of visual information from broader areas, mitigates spatial bias, and thereby enhances the robustness of cross-modal understanding. More cases are available in Appendix D.
@@ 6 Conclusion
This work presents a systematic investigation of spatial bias in LVLMs, showing that inconsistent predictions under spatial shifts originate from unbalanced position encodings in the LLM. To address this issue, we propose Balanced Position Assignment (BaPA), which assigns identical position embeddings to all image tokens. Experiments demonstrate that BaPA improves spatial robustness without retraining and further enhances performance across multimodal benchmarks with lightweight fine-tuning. Further analysis demonstrates that BaPA effectively balances attention distributions, enabling more comprehensive visual understanding.
Ethics Statement
All datasets used in this work are publicly available and widely adopted in the research community. We have manually inspected the data to ensure compliance with the ICLR Code of Ethics. No offensive, harmful, or privacy-sensitive content is involved in our study.
Reproducibility statement
To facilitate reproducibility, we have released the code and part of the data at https://github.com/AAAndy-Zhu/BaPA. Meanwhile, the implementation details of each experiments are provided in Section 3.1, Section 4.1, Section 4.2, Section 5.3, Section 5.4 and Section 5.5, as well as in Appendix C.3. The complete probing dataset and checkpoints of each LVLM will be made publicly available on HuggingFace upon acceptance.
Appendix A Limitations
This paper presents a systematic investigation of spatial bias in LVLMs. Although providing new insights and solutions, our work suffers from two limitations. First, our probing experiments focus primarily on a $3\times 3$ spatial grid, which may not fully capture more fine-grained or irregular spatial variations. Second, due to computational constraints, we limit our evaluation to medium-scale LVLMs and do not include very large models such as Qwen2.5-VL-72B. Investigating whether the identified spatial bias persists or evolves in larger-scale models remains an important direction for future work.
Appendix B Details of Probe Task
For each key image $I_{m}$ and caption $C_{m}$ randomly selected from LAION, we first generate 9 corresponding composite images $\{\bm{I}_{m,n}\}^{8}_{n=0}$ according to the workflow depicted in Figure 6. Then we fed each composite image $\bm{I}_{m,n}$ alongside the question $Q_{m}$ format as following into the LVLMs to test their spatial robustness.
@@ Appendix B Details of Probe Task
Figure 8: Workflow illustration on our analysis experiments.
Appendix C Experiments on Downstream Tasks
C.1 Datasets
@@
CRPE is a benchmark designed to quantitatively evaluate the object recognition and relation comprehension ability of LVLMs. The evaluation is formulated as single-choice questions. Following previous work (Bai et al., 2025), we evaluate LVLMs on relation comprehension ability instead of recognition with CRPE.
@@
C.2 Baselines
CCA
Concentric Causal Attention (CCA) (Xing et al., 2024) is a simple yet effective positional alignment strategy designed for mitigating object hallucination in LVLMs. The core novelty of CCA lies in naturally reducing relative distance between visual and instruction tokens, thereby mitigating the long-term decay effects of RoPE in LVLMs. CCA trains LLaVA from scratch through two stages to adapt the new position embedding, including 1) a pre-training over CC-558K dataset with global batch size of 256 and 2) a instruction tuning with a 665k multi-turn conversation dataset with global batch size of 128. During the experiment, we directly use the checkpoints provided by Xing et al. (2024) for evaluation [^3] https://huggingface.co/xing0047/cca-llava-1.5-7b.
UAC
Uniform Attention Calibration (UAC) (Zhu et al., 2025c) is a training-free method that removes spatial perception bias estimated from a meaningless input by calibrating biased attention, offering a simple yet effective solution with competitive performance. Due to the unavailability of the relevant code and models, we did not compare our method with UAC on general downstream tasks. Instead, we only report its results on the MME benchmark provided by Zhu et al. (2025c).
@@ C.2 Baselines
Gemma3 8 16 32 1.0e-04 2
LLaVA-v1.6 8 16 32 1.0e-04 1
Qwen2.5-VL 8 16 32 1.0e-04 1
LLaVA-v1.5 128 256 128 2.0e-04 1
@@
C.3 Implementation Details
For each LVLM, we randomly select 10,000 samples from the LLaVA-v1.5 instruction-tuning data [^4] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K/blob/main/llava_v1_5_mix665k.json and fine-tune the models with LoRA (Hu et al., 2022) using the LLaMA-Factory library [^5] https://github.com/hiyouga/LLaMA-Factory. The hyperparameters used for training are shown in Table 4. To better adapt LLaVAv1.5 to BaPA, we retrain it from scratch using the same two stages as (Xing et al., 2024), and the hyperparameters are presented in Table 5 All the experiments are finished on 4 A100 GPUs with 80GB memory.
Appendix D Case Study
We further randomly select one example each from HallusionBench, ScienceQA and CRPE for demonstration. The results are shown in Figures 910 and 11.
Figure 9: An example in HallusionBench, where the input image is resized as processed in LLaVA-v1.6.
Figure 10: An example in ScienceQA, where the input image is resized as processed in LLaVA-v1.6.
Figure 11: An example in CRPE, where the input image is resized as processed in LLaVA-v1.6.
Appendix E The Use of Large Language Models
During the preparation of the manuscript, we employed large language models (LLMs) to correct grammatical errors and typos to improve the fluency and readability of the text. All research ideas, experimental designs, analyses, and conclusions are solely developed by the authors.

### v2

@@ 6.4 Analysis of Information Flow
To analyze the effect of AGCI on cross-modal interactions, we present a case study from TextVQA, with the visualization of information flow in Qwen2.5-VL and the AGCI-finetuned version in Figure 4. The baseline Qwen2.5-VL exhibits diffused and weak attention over the image, where the highest attention is incorrectly concentrated on the upper-left region of the image, which is irrelevant to the question, leading to an incorrect prediction. By contrast, after applying AGCI, attention becomes more concentrated on the regions containing the key textual cues, especially for the region containing the word “British” (the brightest patch), which serves as the critical OCR cue for answering the question. By injecting global context into image tokens, AGCI reduces the dominance of position-dependent local cues and provides a more balanced contextual representation, enabling the model to better identify and attend to task-relevant regions.
Figure 5: Effect of $\lambda$ on the Probe Dataset.
6.5 Impact of $\lambda$
To study the influence of the injection strength in AGCI, we vary the injection weight $\lambda$ and evaluate model performance on 9K probe samples from our dataset. As shown in Figure 5, Gemma3n exhibits consistently strong performance across different $\lambda$, achieving high accuracy while maintaining very low variance. As $\lambda$ increases, its overall performance shows a gradual improvement, accompanied by a decreasing trend in variance, suggesting that for robust LVLMs, increasing global context injection within a reasonable range can further enhance cross-modal information integration while preserving stable and reliable spatial reasoning. For LLaVA-v1.6, increasing $\lambda$ leads to a steady improvement in accuracy, indicating that AGCI enables the weaker model to incorporate richer visual information during cross-modal interaction, thereby supporting more accurate understanding and reasoning. However, the variance exhibits a non-monotonic trend, peaking at intermediate values of $\lambda$. We assume that for weaker LVLMs with lower baseline performance, smaller values of $\lambda$ are preferable, as excessive global context injection may overly perturb token representations, leading to increased instability and degraded robustness.
@@ 7 Conclusion
This work presents a systematic investigation of spatial bias in LVLMs, showing that inconsistent predictions under spatial shifts originate from the disruption of global information flow in the LLM. To address this issue, we propose Adaptive Global Context Injection (AGCI), which injects the global visual context into each image token based on semantic similarity. Experiments demonstrate that AGCI not only enhances the spatial robustness but also improves LVLMs’ performance across multimodal benchmarks with lightweight fine-tuning. Further analysis on cross modal information flow emphasizes the importance of incorporating global visual context.
Impact Statements
This paper presents work whose goal is to elucidate and advance the spatial robustness of large vision-language models. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Details of Probe Task
For each key image $I_{m}$ and caption $C_{m}$ randomly selected from LAION, we first generate 9 corresponding composite images $\{\boldsymbol{I}_{m,n}\}^{8}_{n=0}$ according to the workflow depicted in Figure 6. Then we fed each composite image $\boldsymbol{I}_{m,n}$ alongside the question $Q_{m}$ format as following into the LVLMs to test their spatial robustness.
@@ Appendix A Details of Probe Task
Figure 7: Workflow illustration on our analysis experiments.
Appendix B Experiments on Different Image Resolutions
To further explore the relationship between spatial bias and RoPE, we evaluated the performance of Qwen2.5-VL on our probe datasets featuring input images of varying resolutions. The results are shown in Figure 8. It can be found that although higher resolutions introduce a larger number of visual tokens, the overall spatial bias patterns remain largely consistent across resolutions. Notably, we observe that lower-resolution inputs even exhibit more severe bias at positions farther away from the text tokens compared to higher-resolution settings. This counterintuitive behavior suggests that spatial bias does not scale with token distance as would be expected from RoPE-induced long-term decay.
Figure 8: Results on our probe dataset with different resolution of the input image.
@@ Appendix B Experiments on Different Image Resolutions
Appendix C Mutual Information Analysis of AGCI
We state a simple sufficient condition under which AGCI is guaranteed to not reduce the information available for text prediction.
Proposition C.1.
[…]
$I(\tilde{v}_{i}^{\prime},\tilde{g};r)\geq I(\tilde{v}_{i};r)\geq I(\tilde{v}_{i};y).$ (8)
Proof.
Since $y$ is a deterministic function of $r$ (via a fixed output head), the data processing inequality gives
$I(\tilde{v}_{i};y)\leq I(\tilde{v}_{i};r).$ (9)
[…]
$I(\tilde{v}_{i};r)\leq I(\tilde{v}_{i},\tilde{g};r).$ (10)
[…]
$I(\tilde{v}_{i},\tilde{g};r)\leq I(\tilde{v}_{i}^{\prime},\tilde{g};r).$ (11)
Combining these inequalities yields the claim. ∎
Consequently, once the model has access to (an estimate of) the global context $\tilde{g}$, any preferred spatial token can provide at least as much predictive information as the original token, which helps mitigate spatial bias caused by uneven token preference.
Appendix D Experimental Settings
D.1 Downstream Datasets
@@
CRPE is a benchmark designed to quantitatively evaluate the object recognition and relation comprehension ability of LVLMs. The evaluation is formulated as single-choice questions. Following previous work (Bai et al., 2025b), we evaluate LVLMs on relation comprehension ability instead of recognition with CRPE.
@@
TextVQA
TextVQA requires models to read and reason about text in an image to answer questions based on them, containing contains 45,336 questions on 28,408 images. We test each LVLM on the validation set with 5,000 smaples and evaluate the results through the scripts provided by Liu et al. (2023) [^2] https://github.com/haotian-liu/LLaVA/blob/main/scripts/v1_5/eval/textvqa.sh.
POPE
Polling-based Object Probing Evaluation (POPE), a simple yet effective approach for evaluating object hallucination in LVLMs (Li et al., 2023b). It formulates the evaluation as a binary classification task that prompts LVLMs to output “Yes” or “No” and devise three object sampling strategies, including random sampling, popular sampling and adversarial sampling. We evaluate each LVLM with POPE bulit on the validation set of MSCOCO, containing 500 images with 9,000 questions in total.
D.2 Implementation Details
For each LVLM, we randomly select 10,000 samples from the LLaVA-v1.5 instruction-tuning data [^3] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K/blob/main/llava_v1_5_mix665k.json and fine-tune the models with LoRA (Hu et al., 2022) using the LLaMA-Factory library [^4] https://github.com/hiyouga/LLaMA-Factory. The hyperparameters used for training are shown in Table 4. To better adapt LLaVA-v1.5 to AGCI, we retrain it from scratch using the same two stages as Xing et al. (2024), and the hyperparameters are presented in Table 5 All the experiments are finished on 4 NVIDIA H20 GPUs with 96GB memory.
@@ D.2 Implementation Details
Gemma3n-E4B 8 16 32 1.0e-05 2
LLaVA-v1.6-7B 8 16 64 1.0e-04 1
InternVL3-8B 8 16 16 1.0e-04 2
Qwen3-VL-8B 8 16 64 1.0e-04 1
Qwen2.5-VL-7B 8 16 64 1.0e-04 1
Qwen2.5-VL-32B 8 16 64 1.0e-06 1
Qwen2.5-VL-72B 8 16 64 1.0e-04 1
@@
Table 6: Results on POPE dataset. The best results are bolded
Models Adversarial Popular Random Average
Acc F1 Acc F1 Acc F1 Acc F1
Gemma3n 86.77 85.61 87.90 86.69 88.93 87.69 87.87 86.66
Gemma3n-AGCI 87.07 86.11 88.37 87.32 89.40 88.32 88.28 87.25
LLaVA-v1.6 87.10 86.55 88.70 88.02 90.80 90.02 88.87 88.20
LLaVA-v1.6-AGCI 87.17 86.79 89.23 88.69 91.40 90.75 89.27 88.74
InternVL3 86.43 85.43 88.00 86.90 89.20 88.05 87.88 86.79
InternVL3-AGCI 86.57 85.70 88.13 87.16 89.57 88.53 88.09 87.13
Qwen3-VL-8B 82.89 83.51 84.10 84.49 86.54 86.85 84.51 84.84
Qwen3-VL-AGCI 84.67 84.96 86.57 86.58 89.63 89.32 86.96 86.95
Qwen2.5-VL-7B 87.77 88.12 90.67 90.67 94.13 93.93 90.86 90.91
Qwen2.5-VL-AGCI 89.00 88.88 90.73 90.47 93.17 92.79 90.97 90.72
LLaVA-v1.5 78.96 77.57 81.88 80.06 83.29 81.33 81.38 79.65
LLaVA-v1.5-VCD 80.88 81.33 85.38 85.06 87.73 87.16 84.66 84.52
LLaVA-v1.5-RLHF 82.60 80.88 83.90 82.05 85.90 83.92 84.13 82.28
LLaVA-v1.5-CCA 85.67 84.42 86.87 85.54 88.03 86.65 86.86 85.54
LLaVA-v1.5-AGCI 85.50 84.51 86.87 85.76 88.57 87.37 86.98 85.88
Appendix E Case Study
We further randomly select one example CRPE for demonstration. The results are shown in Figures 11.
Figure 11: An example from CRPE benchmark.